Give an example of seperable Hausdroff space which has a non seperable subspace


1
Expert's answer
2021-07-21T17:39:02-0400

Consider "\mathbb{R}" with the lower-limit topology, i.e. the topology generated by the base "\{[a,b)|a,b\in \mathbb{R} \text{ and } a<b\}". If we generalize this topology on "\mathbb{R}^2", the topology we'll get is finer than the usual topology on "\mathbb{R}^2". So it is Hausdorff and separable.

"\{(x,-x)|x \in \mathbb{R}\}" is an uncountable discrete subspace of this topological space.